# Subgraph-Aware Training of Text-based Methods for Knowledge Graph Completion

### v3

@@ Abstract
Fine-tuning pre-trained language models (PLMs) has recently shown a potential to improve knowledge graph completion (KGC). However, most PLM-based methods encode only textual information, neglecting various topological structures of knowledge graphs (KGs). In this paper, we empirically validate the significant relations between the structural properties of KGs and the performance of the PLM-based methods. To leverage the structural knowledge, we propose a Subgraph-Aware Training framework for KGC (SATKGC) that combines (i) subgraph-aware mini-batching to encourage hard negative sampling, and (ii) a new contrastive learning method to focus more on harder entities and harder negative triples in terms of the structural properties. To the best of our knowledge, this is the first study to comprehensively incorporate the structural inductive bias of the subgraphs into fine-tuning PLMs. Extensive experiments on four KGC benchmarks demonstrate the superiority of SATKGC. Our code is available.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
@@
Factual sentences, e.g., Leonardo da Vinci painted Mona Lisa, can be represented as entities, and relations between the entities. Knowledge graphs treat the entities (e.g., Leonardo da Vinci and Mona Lisa) as nodes, and the relations (e.g., painted) as edges. Each edge and its endpoints are denoted as a triple ($h,r,t$) where $h,r$ and $t$ are a head entity, a relation, and a tail entity respectively. Since KGs can represent complex relations between entities, they serve as key components for knowledge-intensive applications Ji et al. (2021); Xiong et al. (2017); He et al. (2017); Zhang et al. (2016); Liu et al. (2021).
@@
Despite their applicability, real-world KGs miss factual relations, which can be inferred from existing facts in the KGs. Hence, the task of knowledge graph completion (KGC) has become an active research topic Ji et al. (2021). Given an incomplete triple $(h,r,?)$, this task is to predict the correct tail $t$. A true triple $(h,r,t)$ in KG and a false triple $(h,r,\hat{t})$ which does not exist in KG are called positive and negative, respectively. A negative triple difficult for a KGC method to distinguish from its corresponding positive triple is regarded as a hard negative triple.
@@
Existing KGC methods are categorized into two approaches. An embedding-based approach learns embeddings of entities in continuous vector spaces, but ignores contextualized text information in KGs, thus being inapplicable to entities and relations unseen in training (Bordes et al., 2013; Balazevic et al., 2019). A text-based approach, based on pretrained language models (PLMs), learns textual representations of KGs, but suffers from a lack of structural knowledge in KGs (Wang et al., 2022a).
@@
Meanwhile, contrastive learning has become a key component of representation learning Kalantidis et al. (2020); Robinson et al. (2021); Gao et al. (2021); Wang et al. (2022b), but an important aspect of contrastive learning, i.e., the effect of hard negatives, has so far been underexplored in KGC. In this paper, we empirically validate significant relationships between the structural properties of KGs and the performance of the PLM-based KGC methods.
To address the aforementioned longstanding limitations of two KGC approaches by utilizing the above relationships, we hypothesize that incorporating the structural inductive bias of KGs into both sampling hard negatives and fine-tuning PLMs leads to a major breakthrough in learning comprehensive representations of KGs. To this end, we propose a Subgraph-Aware Training framework for KGC (SATKGC), which (i) samples subgraphs of a KG to treat triples of each subgraph as a mini-batch to encourage hard negative sampling, and (ii) fine-tunes a PLM via contrastive learning which focuses more on structurally harder entities and structurally harder negative triples induced by topological bias in the KG. To sum up, we make four contributions.
@@
  • We propose a subgraph-aware training strategy for PLM-based KGC methods, which is effective in sampling in-batch hard negatives.
  • We propose a novel contrastive learning method that gives different importances to positive and negative triples based on the structural properties of KGs.
@@
  • We conduct extensive experiments on four KGC benchmarks to demonstrate the superiority of SATKGC over existing KGC methods.
2 Related Work
@@
An embedding-based approach maps complex and structured knowledge into low-dimensional spaces. This approach computes the plausibility of a triple using translational scoring functions on the embeddings of the triple’s head, relation, and tail Bordes et al. (2013); Sun et al. (2019); Zhang et al. (2022); Ge et al. (2023), e.g., $h+r\approx t$, or semantic matching functions which match latent semantics of entities and relations Nickel et al. (2011); Yang et al. (2015); Trouillon et al. (2016); Balazevic et al. (2019). This approach exploits the spatial relations of the embeddings, but cannot make use of texts in KGs, i.e., the source of semantic relations.
@@
In contrast, a text-based approach learns contextualized representations of the textual contents (e.g., names and descriptions) of entities and relations by leveraging PLMs Yao et al. (2019); Wang et al. (2021b); Xie et al. (2016); Daza et al. (2021); Kim et al. (2020); Wang et al. (2022a); Chen et al. (2022); Zhang et al. (2023a). More recently, there has been a significant increase in the adoption of large language models (LLMs) in KGC Zhang et al. (2023c). However, these PLM-based models are often oblivious to structural inductive bias in KGs.
@@
A few attempts have been made to utilize the above two approaches at once. StAR Wang et al. (2021d) proposes an ensemble model incorporating an output of a Transformer encoder Vaswani et al. (2017) with a triple score produced by RotatE Sun et al. (2019). CSProm-KG Chen et al. (2023) trains KG embeddings through the soft prompt for a PLM. Nevertheless, the integration of structural and textual information in a KG in training has not yet been fully realized.
Contrastive learning, shown to be effective in various fields Wu et al. (2018); Haque et al. (2022); Fang et al. (2020); Zhang et al. (2023b); Gao et al. (2021), has recently emerged as a promising approach in the context of KGC Wang et al. (2022a); Yang et al. (2020); Qiao et al. (2023). Despite a critical aspect of contrastive learning, the effect of hard negatives has been overlooked in KGC.
@@
3 Motivation
To demonstrate the limitations of text-based methods exhibiting competitive performance such as SimKGC (Wang et al., 2022a) and StAR (Wang et al., 2021a), we investigate the characteristics of false positive (FP) triples which are ranked by these models higher than corresponding true triples, on two widely-used datasets WN18RR and FB15k-237. Our analysis draws two conclusions.
Figure 1: False positive (FP) ratio against the distance (i.e., length of the shortest path) between head and tail of a FP triple in KG across different text-based methods.
Figure 2: False positive (FP) ratio against the degree of head for a FP triple across different text-based methods.
First, the closer the tail and head of a false triple are to each other in the KG, the more likely the false triple is to be ranked higher than the corresponding true triple. Figure 1 illustrates the distribution of distance, i.e., the length of the shortest path, between the head and tail of a FP triple, where $y$-axis represents the FP ratio111(the number of FPs with a specific head-to-tail distance) / (the number of pairs of entities in the KG with the same distance). for each distance. For StAR and SimKGC, the FP ratio dramatically grows as the distance decreases (see green and red bars in Figure 1). These findings highlight the importance of considering the proximity of two entities of a negative triple in KGs for text-based methods to distinguish a positive triple from the negative.
Figure 3: Overview of the proposed training framework, which consists of: (i) sampling subgraphs from KG (before training); (ii) selecting a subgraph; (iii) fetching a mini-batch of triples in the subgraph; (iv) calculating a similarity between embeddings of every (head, relation) pair and every tail in the mini-batch; (v) contrastive learning via infoNCE loss incorporated with two structure-aware factors.
Second, we discover that the higher the degree of the head in a false triple is, the more likely the false triple is to be ranked higher than the corresponding true triple. Figure 2 illustrates the distribution of the degree of heads of FP triples. The FPs are sorted in ascending order of the degrees, and then they are divided into five groups such that each group contains an equal number of distinct degrees. The $y$-axis represents the FP ratio222(the average number of FPs whose head’s degree falls into each group) / (the number of entities in the KG whose degree falls into each group) in each degree group. The FP ratio for StAR and SimKGC increases as the degree of the head grows (see green and red bars in Figure 2). This indicates that the existing text-based methods have difficulty in predicting correct tails for missing triples with the high-degree heads. Hence, the degree can be taken into account to enhance the performance of text-based methods.
Our proposed framework (dubbed SATKGC) tackles the above two phenomena [^3] The trends in Figures 1 and 2 are also confirmed on Wikidata5M and NELL-995., thereby significantly reducing the FPs for all distances and all degree groups compared to the existing methods (see blue bars in Figures 1 and 2).
4 Method
We propose a novel training framework for KGC that captures the structural inductive bias of the KG, based on the aforementioned observations.
@@ 4 Method
Figure 3 illustrates the overview of our framework. First, for every triple, a subgraph is extracted around that triple from the KG, performed before training (Section 4.1). During training, we keep track of the number of visits for every triple. For each iteration, a subgraph is selected based on that number, and then all forward and inverse triples in the subgraph are fetched as a mini-batch $B$ to a model (Section 4.2). We adopt the bi-encoder architecture Wang et al. (2022b) as a backbone, which uses pre-trained BERT (Devlin et al., 2018) as encoders. Specifically, $\textit{Encoder}_{hr}$ and $\textit{Encoder}_{t}$ take the text, i.e., name and description, of $(h,r)$ and $t$ as input, and produce their embeddings $x_{hr}$ and $x_{t}$ respectively. We then calculate the cosine similarity between $x_{hr}$ and $x_{t}$ for every $(h,r)$ and $t$ in the mini-batch, and perform contrastive learning based on two structure-aware factors (Section 4.3). The model inference is described in Appendix A.
@@
Figure 4: Example of BRWR-based subgraph sampling; (a) probability of selecting the start entity $s$ between $h$ and $t$ of a center triple, where $t$ with a lower degree is more likely to be selected as $s$ than $h$; (b) probability of selecting a neighbor of current entity $u$. A random walker is more likely to move to $v_{1}$ than to $v_{2}$ whose degree is larger than $v_{1}$.
4.1 Random-Walk Based Subgraph Sampling
@@
First, we aim to extract subgraphs from the KG to treat all the triples in the extracted subgraph as a mini-batch for training. For each triple in the KG, we perform BRWR starting from that triple called a center triple, and the triples visited by BRWR composes an extracted subgraph as follows: (i) for each center triple, either head $h$ or tail $t$ of the center triple is selected as the start entity $s$ based on an inverse degree distribution of $h$ and $t$, i.e., $\frac{|N(x)|^{-1}}{|N(h)|^{-1}+|N(t)|^{-1}}$, where $x\in\{h,t\}$ and $N(x)$ denotes a set of $x$’s neighbors; (ii) next, we perform BRWR from $s$ until we sample $M$ triples where $M$ is a predefined maximum number (e.g., 10,000). For each iteration in BRWR, a random walker moves from the current node to either $s$ with a probability of $p_{r}$ or one of the neighbors of the current node with a probability of $1-p_{r}$. We define the probability of selecting one of $u$’s neighbors $v\in N(u)$ as $p_{v}=\frac{|N(v)|^{-1}}{\sum_{v\in N(u)}|N(v)|^{-1}}$, which is a normalized inverse degree distribution of the neighbors. Figures 4(a) and 4(b) show the running example of step (i) and an iteration of step (ii).
Performed before the model training, this subgraph sampling algorithm can extract many distinct entities due to the inverse degree distribution, which will be validated in Section 5.5.
4.2 Subgraph as a Mini-Batch

### v5

# Subgraph-Aware Training of Language Models for Knowledge Graph Completion Using Structure-Aware Contrastive Learning

@@ Abstract.
Fine-tuning pre-trained language models (PLMs) has recently shown a potential to improve knowledge graph completion (KGC). However, most PLM-based methods focus solely on encoding textual information, neglecting the long-tailed nature of knowledge graphs and their various topological structures, e.g., subgraphs, shortest paths, and degrees. We claim that this is a major obstacle to achieving higher accuracy of PLMs for KGC. To this end, we propose a Subgraph-Aware Training framework for KGC (SATKGC) with two ideas: (i) subgraph-aware mini-batching to encourage hard negative sampling and to mitigate an imbalance in the frequency of entity occurrences during training, and (ii) new contrastive learning to focus more on harder in-batch negative triples and harder positive triples in terms of the structural properties of the knowledge graph. To the best of our knowledge, this is the first study to comprehensively incorporate the structural inductive bias of the knowledge graph into fine-tuning PLMs. Extensive experiments on three KGC benchmarks demonstrate the superiority of SATKGC. Our code is available. [^1] https://github.com/meaningful96/SATKGC
Knowledge Graph Completion, Contrastive Learning, Hard Negative Sampling
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714946isbn: 979-8-4007-1274-6/25/04ccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
Factual sentences, e.g., Leonardo da Vinci painted Mona Lisa, can be represented as entities, and relations between the entities. Knowledge graphs treat the entities (e.g., Leonardo da Vinci, Mona Lisa) as nodes, and the relations (e.g., painted) as edges. Each edge and its endpoints are denoted as a triple ($h,r,t$), where $h,r$, and $t$ are a head entity, a relation, and a tail entity respectively. Since KGs can represent complex relations between entities, they serve as key components for knowledge-intensive applications (Ji et al., 2021; Xiong et al., 2017; He et al., 2017; Zhang et al., 2016; Liu et al., 2021).
@@
Despite their applicability, factual relations may be missing in incomplete real-world KGs, and these relations can be inferred from existing facts in the KGs. Hence, the task of knowledge graph completion (KGC) has become an active research topic (Ji et al., 2021). Given an incomplete triple $(h,r,?)$, this task is to predict the correct tail $t$. A true triple $(h,r,t)$ in a KG and a false triple $(h,r,\hat{t})$ which does not exist in the KG are called positive and negative, respectively.
A negative triple difficult for a KGC method to distinguish from its corresponding positive triple is regarded as a hard negative triple, i.e., a false triple $(h,r,\hat{t})$ whose tail $\hat{t}$ is close to tail $t$ of the positive triple $(h,r,t)$ in the embedding space.
Figure 1. False positive (FP) ratio against the distance (i.e., length of the shortest path) between head and tail of a FP triple in KG across different text-based methods.
Figure 2. False positive (FP) ratio against the degree of tail for a FP triple across different text-based methods.
@@
Existing KGC methods are categorized into two approaches. An embedding-based approach learns embeddings of entities in continuous vector spaces, but ignores contextualized text information in KGs, thus being inapplicable to entities and relations unseen in training (Bordes et al., 2013; Balazevic et al., 2019). A text-based approach, based on pretrained language models (PLMs), learns textual representations of KGs, but suffers from a lack of structural knowledge inherent in KGs (Wang et al., 2022). Moreover, most methods in this approach fail to account for the long-tailed distribution of entities in KGs, though they can perform KGC even in the inductive setting [^2] In the inductive setting for KGC, entities in the test set never appear in the training set..
@@
Meanwhile, contrastive learning has become a key component of representation learning (Kalantidis et al., 2020; Robinson et al., 2021; Gao et al., 2021; Wang et al., 2022), but an important aspect of contrastive learning, i.e., the effect of hard negatives, has so far been underexplored in KGC.
In this paper, we empirically validate significant relationships between the structural inductive bias of KGs and the performance of the PLM-based KGC methods. To demonstrate the limitations of language models exhibiting competitive performance such as SimKGC (Wang et al., 2022) and StAR (Wang et al., 2021b), we investigate the characteristics of false positive (FP) triples, i.e., false triples ranked higher than a true triple by the models, on two well-known datasets WN18RR and FB15k-237. Our analysis draws two observations.
First, the closer the tail and head of a false triple are to each other in the KG, the more likely the false triple is to be ranked higher than the corresponding true triple. Figure 1 illustrates the distribution of distance, i.e., the length of the shortest path, between the head and tail of a FP triple, where $y$-axis represents the FP ratio333(the number of FPs with a specific head-to-tail distance) / (the number of pairs of entities in the KG with the same distance) for each distance. For StAR and SimKGC, the FP ratio dramatically grows as the distance decreases (see green and red bars in Figure 1). These findings highlight the importance of considering the proximity between two entities in a triple to distinguish between true and false triples.
Second, the higher the degree of the tail in a false triple is, the more likely that triple is to be ranked higher than the corresponding true triple. Figure 2 illustrates the distribution for the degree of tails of FP triples. They are sorted in ascending order of their degrees, and then are divided into five groups such that each group contains an equal number of distinct degrees. The $y$-axis represents the FP ratio444(the average number of FPs whose tail’s degree falls into each group) / (the number of entities in the KG whose degree falls into each group) in each degree group. The FP ratio for StAR and SimKGC increases as the degree of the tail grows (see green and red bars in Figure 2). This indicates that the existing text-based methods have difficulty in predicting correct tails for missing triples with the high-degree tails. Hence, the degree can be taken into account to enhance the performance of language models.
In this paper, we tackle the above two phenomena [^5] The trends in Figures 1 and 2 are also confirmed on Wikidata5M., thereby significantly reducing the FPs compared to the existing methods (see blue bars in Figures 1 and 2). For this, we hypothesize that incorporating the structural inductive bias of KGs into hard negative sampling and fine-tuning PLMs leads to a major breakthrough in learning comprehensive representations of KGs.
To this end, we propose a subgraph-aware PLM training framework for KGC (SATKGC) with three novel techniques: (i) we sample subgraphs of the KG, and treat triples of each subgraph as a mini-batch to encourage hard negative sampling and to alleviate the negative effects of long-tailed distribution of the entity frequencies during training; (ii) we fine-tune PLMs via a novel contrastive learning method that focuses more on harder negative triples, i.e., negative triples whose tails can be reached from their heads by a smaller number of hops in the KG; and (iii) we propose balanced mini-batch loss that mitigates the gap between the long-tailed distribution of the training set and a nearly uniform distribution of each mini-batch. To sum up, we make three contributions.
@@
  • We propose a new training strategy for PLMs, which not only effectively samples hard negatives from a subgraph but also visits all entities in the KG nearly equally in training, as opposed to the existing PLM-based KGC methods. Based on the structural properties of KGs, our contrastive learning enables PLMs to pay attention to difficult negative triples in KGs, and our mini-batch training eliminates the discrepancy between the distributions of a training set and a mini-batch.
@@
  • We conduct extensive experiments on three KGC benchmarks to demonstrate the superiority of SATKGC over existing KGC methods.
2. RELATED WORK
@@
An embedding-based approach maps complex and structured knowledge into low-dimensional spaces. This approach computes the plausibility of a triple using translational scoring functions on the embeddings of the triple’s head, relation, and tail (Bordes et al., 2013; Sun et al., 2019; Zhang et al., 2022; Ge et al., 2023), e.g., $h+r\approx t$, or semantic matching functions which match latent semantics of entities and relations (Nickel et al., 2011; Yang et al., 2015; Trouillon et al., 2016; Balazevic et al., 2019). KG-Mixup (Shomer et al., 2023) proposes to create synthetic triples for a triple with a low-degree tail. UniGE (Liu et al., 2024) utilizes both hierarchical and non-hierarchical structures in KGs. The embedding-based approach exploits the spatial relations of the embeddings, but cannot make use of texts in KGs, i.e., the source of semantic relations.
@@
In contrast, a text-based approach learns contextualized representations of the textual contents of entities and relations by leveraging PLMs (Yao et al., 2019; Wang et al., 2021a; Xie et al., 2016; Daza et al., 2021; Kim et al., 2020; Wang et al., 2022; Chen et al., 2022; Wang et al., 2023). Recently, a sequence-to-sequence model for KGC (Yu and Yang, 2023) highlights the growing trend of leveraging natural language generation models. With the rise of large language models (LLMs), their application in KGs has significantly increased (Zhang et al., 2023a). Despite these advancements, PLMs often lack awareness of the structural inductive bias inherent in KGs.
@@
A few attempts have been made to utilize the above two approaches at once. StAR (Wang et al., 2021b) proposes an ensemble model incorporating an output of a Transformer encoder (Vaswani et al., 2017) with a triple score produced by RotatE (Sun et al., 2019). CSProm-KG (Chen et al., 2023) trains KG embeddings through the soft prompt for a PLM. Nevertheless, the integration of structural and textual information in a KG in training has not yet been fully realized.
Contrastive learning, shown to be effective in various fields (Wu et al., 2018; Haque et al., 2022; Fang et al., 2020; Zhang et al., 2023b; Gao et al., 2021), has recently emerged as a promising approach in the context of KGC (Wang et al., 2022; Yang et al., 2020; Qiao et al., 2023). KRACL (Tan et al., 2023) introduces contrastive learning into an embedding-based method, particularly through the use of a graph neural network (GNN), while HaSa (Zhang et al., 2023c) applies contrastive learning to a PLM. HaSa aims to sample hard negatives which are unlikely to be false negatives by selecting tails of a negative triple based on the number of 1- and 2-hop neighbors of its head entity, i.e., the greater the number, the lower the probability that the tail is included in the false negative. In contrast, our contrastive learning method estimates the difficulty of negative triples by utilizing various length of the shortest path between their head and tail, while simultaneously mitigating the adverse effects caused by the long-tailed distribution in KGs. The long-tailed distribution problem, which reflects class imbalance in data, has been extensively studied in other domains, particularly computer vision, where numerous efforts have been made to address this challenge (Tan et al., 2020; Jamal et al., 2020; Li et al., 2020; Zhu et al., 2022; Ren et al., 2020).
@@
Figure 3. Overview of the proposed training framework, which consists of: (i) Random-walk Based Subgraph Sampling (before training); (ii) Subgraph as a Mini-batch; (iii) Proximity-aware Contrastive Learning; (iv) Frequency-aware Mini-batch Training.
3. PRELIMINARY
Let $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$ denote a KG in which $\mathcal{E}$ and $\mathcal{R}$ represent a set of entities and relations, respectively, and $\mathcal{T}=\{(h,r,t)|h,t\in\mathcal{E},r\in\mathcal{R}\}$ is a set of triples where $h$, $r$, and $t$ are a head entity, a relation, and a tail entity, respectively.
Problem Definition.Given a head $h$ and a relation $r$, the task of knowledge graph completion (KGC) aims to find the most accurate tail $t$ such that a new triple $(h,r,t)$ is plausible in $\mathcal{G}$.
InfoNCE Loss. InfoNCE loss (Oord et al., 2018) has been widely employed to learn the representations for audio, images, natural language, and even for KGs (Zhang et al., 2023c; Wang et al., 2022). Given a set $X=\{t_{1},t_{2},...,t_{n}\}$ of $n$ tails containing one positive sample from $p(t_{j}|h,r)$ and $n-1$ negative samples from the proposal distribution $p(t_{j})$, InfoNCE loss (Wang et al., 2022) for KGC is defined as:
$\mathcal{L}_{X}=\mathbb{E}_{X}\left[-\log\frac{\text{exp}(\phi(h,r,t))}{\sum\limits_{t_{j}\in X}\text{exp}(\phi(h,r,t_{j}))}\right]\end{aligned}$
$\phi(h,r,t)=\text{cos}(\mathbf{x}_{hr},\mathbf{x}_{t})\end{aligned}$
where a scoring function $\phi(h,r,t)$ is defined as the cosine similarity between two representations $\mathbf{x}_{hr}$ for $h$ and $t$, and $\mathbf{x}_{t}$ for $t$.
4. METHOD
@@
In this section, we describe a novel PLM fine-tuning framework for KGC. Figure 3 illustrates the overview of our framework for a given KG. First, for every triple, a subgraph is extracted around that triple from the KG before training (Section 4.1). During training, we keep track of the number of visits for every triple. For each iteration, a subgraph is selected based on that number, and then all forward and inverse triples in the subgraph are fetched as a mini-batch $\mathcal{B}$ to the language model (Section 4.2). We adopt the bi-encoder architecture (Wang et al., 2022) with two pre-trained MPNets (Song et al., [n. d.]) as a backbone. Specifically, $\textsf{Encoder}_{hr}$ and $\textsf{Encoder}_{t}$ take the text, i.e., name and description, of $(h,r)$ and $t$ as input, and produce their embeddings $\mathbf{x}_{hr}$ and $\mathbf{x}_{t}$ respectively. We then calculate the cosine similarity between $\mathbf{x}_{hr}$ and $\mathbf{x}_{t}$ for every $(h,r)$ and $t$ in the mini-batch, and perform new contrastive learning equipped with two topology-aware factors (Section 4.3 and 4.4). The details of the input format are provided in Appendix D, and the model inference is described in Appendix A.
4.1. Random-walk Based Subgraph Sampling
@@
Figure 4. Example of BRWR-based subgraph sampling; (a) probability of selecting start entity $s$ between $h$ and $t$ of a center triple, where $t$ with a lower degree is more likely to be $s$ than $h$; (b) probability of selecting a neighbor of current entity $u$. A random walker is more likely to move to $v_{1}$ than to $v_{2}$ with its degree larger than $v_{1}$.
@@
We aim to extract subgraphs from the KG to treat all the triples in the extracted subgraph as a mini-batch for training. For each triple in the KG, therefore, we perform BRWR starting from that triple called a center triple, and the triples visited by BRWR compose the subgraph before training as follows: (i) either head $h$ or tail $t$ of the center triple is selected as the start entity $s$ based on an inverse degree distribution of $h$ and $t$, i.e., $\frac{|N(v)|^{-1}}{|N(h)|^{-1}+|N(t)|^{-1}}$, where $v\in\{h,t\}$ and $N(v)$ denotes a set of $v$’s neighbors; (ii) next, we perform BRWR from $s$ until we sample $M$ triples where $M$ is a predefined maximum number (e.g., 10,000). For each iteration in BRWR, a random walker moves from the current node to either $s$ with a probability of $p_{r}$ or one of the neighbors of the current node with a probability of $1-p_{r}$. We define the probability of selecting one of $u$’s neighbors $v\in N(u)$ as $p_{v}=\frac{|N(v)|^{-1}}{\sum_{v_{j}\in N(u)}|N(v_{j})|^{-1}}$, which is a normalized inverse degree distribution of the neighbors. Figures 4(a) and 4(b) show the running example of step (i) and an iteration of step (ii).
In this manner, giving more chances for an entity with a lower degree to engage in training alleviates the skewed frequency of entity occurrences throughout training, which in turn enhances the KGC performance. This claim will be validated in Section 5.5.
4.2. Subgraph as a Mini-batch